Let h(x)= (g o f)(x)= x^2 / x^2 + 1
Which of the following could be a possible decomposition of h(x)?
a. f(x)= x; g(x)= x / x + 1
b. f(x)= x + 1; g(x)= 1 / x^2
c. f(x)= x^2; g(x)= x / x + 1
d. f(x)= x + 1; g(x)= x^2
Please help me. It is multiple choice, only one answer!

Answer by rothauserc(4718)

for x not = to 0, we have
Let h(x)= (g o f)(x)= (x^2 / x^2) + 1 = 2
correct answer is a.
(g o f)(x)= (x / x) + 1 = 2